### v1

@@ 1 Introduction
The inference problem considered here departs from the study of static systems and instead addresses systems that can evolve over time. We assume that inference is conducted independently over non-overlapping epochs, each treated as locally stationary so that asymptotic estimation theory applies within each epoch. The sample size denotes the number of observations allocated to a single epoch and may vary between epochs to reflect changes in operating conditions. The variance may likewise change, reflecting evolution of the underlying system. In this way, non-stationarity is accommodated through shifts in thermodynamic coordinates between epochs.
@@
The main contributions of the paper are as follows: (1) a thermodynamic state-space formulation of asymptotic inference in $(m,\sigma^{2})$ space, with an explicit uncertainty state function; (2) a reversed second-law inequality for mean inference, together with an integrating factor that yields a first-law–type balance relation; (3) optimal information-gain paths and efficiency bounds; and (4) a unification of familiar entropy–information identities, including the Gaussian-limit form of the de Bruijn and I–MMSE results, as coordinate projections of a single underlying structure.
@@ 3 Inference in sensory transduction and metrology
Three additional empirical assumptions are required to complete the description. First, the variance of stimulus fluctuations $\sigma^{2}(\mu)$ is assumed to be a continuous, non-decreasing function of the mean stimulus intensity $\mu$ (c.f. fluctuation scaling [4]). Second, the sample size evolves dynamically, relaxing monotonically toward a unique equilibrium or optimal value $m_{eq}$. Third, since the variance increases monotonically with $\mu$, it is expected that the optimal sample size itself depends monotonically on stimulus magnitude $m_{eq}(\mu)$ although not necessarily with the same functional form as $\sigma^{2}(\mu)$. In what follows, sample size $m$ is treated as a continuous variable, with the approximation justified in the asymptotic regime.
@@
Sensory systems are not unique in exhibiting this statistical structure. A similar description arises whenever an unknown parameter is inferred from repeated, independent measurements. In metrology, one estimates a parameter $\theta$ of a distribution $p(x\mid\theta)$ from a set of finite samples $x_{1},\ldots,x_{m}$. In the asymptotic regime, efficient estimators become Gaussian with variance proportional to $1/m$, Fisher information becomes additive in $m$, and the estimator entropy admits a closed-form expression identical to (2) with $\sigma^{2}(\theta)=1/I(\theta)$. The inferential dynamics therefore run in the same direction as in sensory systems: increasing sample size reduces uncertainty. What metrology lacks, without further specification, is a dynamical sampling model (1), and the identification of uncertainty with a directly observable response. Moreover, without the analogue of $F=kH$, the thermodynamic structure describes limits on inferential efficiency rather than experimentally measurable results. Nevertheless, the underlying structure remains the same. In this sense, sensory transduction and measurement theory occupy the same position within a single inferential state space.
Next we turn to developing this framework further by quantifying entropy production in inference and examining the role of the second law.
@@ 4 Foundations for a thermodynamic framework
where $\delta H_{\mathrm{flux}}$ is the reversible entropy change due to externally driven changes of the stimulus or the system, and $\delta H_{\mathrm{relax}}$ is the internal response associated with sampling. Defining
$d\mathscr{I}=-\,\delta H_{\mathrm{relax}},$ (6)
we refer to this quantity simply as information. Under the conditions discussed in Appendix 2, the theorem there gives the cyclic inequality
@@ 5 Identifying an integrating factor and a first law
The term $\partial H/\partial m$ represents relaxation and corresponds to “irreversible” entropy production, whereas $\partial H/\partial\sigma^{2}$ represents entropy flux and defines the “reversible” contribution:
$dH_{rev}=\frac{d\sigma^{2}}{2(\sigma^{2}+m\sigma_{R}^{2})}.$ (10)
Next, we introduce the new state variable
@@
so that the entropy change takes the Clausius-like form of $dH_{rev}=\Theta^{-1}d\sigma^{2}$. Here the state variable $\Theta$ plays a role analogous to temperature in thermal physics: it acts as an integrating factor for the reversible entropy flux. As will be discussed later, at fixed large $m$, $\Theta$ is proportional to the inverse minimum mean square error (MMSE) for an averaged signal corrupted by additive Gaussian noise. An increase in variance produces a larger change in entropy when $m$ is small than when $m$ is large. For this reason, we refer to $\Theta$ as the uncertainty susceptibility.
It is important to emphasize that the two coordinates $m$ and $\sigma^{2}$ are treated independently and interpreted operationally. The quantity $m$ denotes the number of observations assigned to a single inference interval, with intervals taken to be statistically independent and non-overlapping. $\sigma^{2}$ can change across different epochs reflecting changes undergone in the system. Asymptotic relations apply within each interval, with no information carried between intervals. Accordingly, variations in $(m,\sigma^{2})$ represents shifts in operating conditions.
Interpreting variance in (10) as the quasi-heat is natural: variance is positive and quadratic, and often corresponds directly to signal power. Next, we can write the equivalent first law. From the differential balance equation (9), and introducing $\Theta$, we get
@@
The first law can also be illustrated with some full examples. The quasi-work term or sampling work $\int_{m_{a}}^{m_{b}}(\sigma^{2}/m)dm$ can be calculated easily when the variance is constant giving $\sigma^{2}\log(m_{b}/m_{a})$. The term $(\sigma^{2}/m)\,dm$ can be interpreted as the incremental variance expenditure required to increase the effective sample size by $dm$. This expression has the units of variance and reflects both the precision as well as the benefit or effort from changing the sample size from $m_{a}$ to $m_{b}$. The term reflects the fact that precision falling off as $1/m$ with a decreasing rate of return when sample size is increased. When $m_{a}=m_{b}$, the sampling work equals zero. In general, the integral is path dependent in $(m,\sigma^{2})$ state space as $(\sigma^{2}/m)\,dm$ is not an exact differential, and its value is non-zero when evaluated over a cycle. Consider a system with $\sigma_{a}^{2}$ sampled with sample size $m_{a}$. At $t=0$, the system is perturbed such that the variance increases to $\sigma_{b}^{2}\geq\sigma_{a}^{2}$ with the sample size increased to $m_{b}$. At a later time, the variance is returned to $\sigma_{a}^{2}$ and the sample size is returned to $m_{a}$. The sampling work over this cycle is
@@
$=(\sigma_{b}^{2}-\sigma_{a}^{2})\log(m_{b}/m_{a}),$ (20)
which can be interpreted geometrically as a weighted area enclosed in $(m,\sigma^{2})$ space.
@@ 6 Other thermodynamic relationships, and a third law
$\left(\frac{\partial H}{\partial\sigma^{2}}\right)_{m}$ $=\frac{1}{\Theta},\quad$ $\left(\frac{\partial H}{\partial m}\right)_{\sigma^{2}}$ $=-\frac{\sigma^{2}}{\Theta m},$ (22)
@@ 7.1 Information gain
$\Delta\mathscr{I}=-\displaystyle\int_{m_{a}}^{m_{b}}\left(\frac{\partial H}{\partial m}\right)\,dm=\displaystyle\int_{m_{a}}^{m_{b}}\frac{1}{\Theta}\frac{\sigma^{2}}{m}\,dm,$ (23)
@@ 7.2 Information efficiency
The estimator variance is $\sigma^{2}/m$, and so increasing the sample size reduces the effective variance. In this sampling-driven setting, (29) corresponds to the Gaussian form of de Bruijn’s identity [3], with $m$ playing the role of a time or noise parameter through the ratio $\sigma^{2}/m$. Alternatively, if variance is treated as the independent control variable, $\Theta$ is related to the minimum mean-square error (MMSE) of a Gaussian channel. For an additive Gaussian channel with signal-to-noise ratio $\mathrm{SNR}=\sigma^{2}/m\sigma_{R}^{2}$, one obtains
@@
In this representation, (29) reduces to the I–MMSE relation [6]. The appearance of both de Bruijn’s identity and the I–MMSE relation reflects the Gaussian structure underlying the entropy function. More importantly, both arise as distinct coordinate projections of a single underlying thermodynamic structure. This unification is encoded by the state function $\Theta=2(\sigma^{2}+m\sigma_{R}^{2})$, which governs joint sampling- and variance-driven information transduction.
@@
where normalization by the input signal variance $\sigma^{2}/m$ defines a dimensionless quantity $\eta$ characterizing local information efficiency. The efficiency obeys $0\leq\eta\leq 1$, with $\eta=1$ attained at the noise floor. The existence of a temperature-like variable $\Theta$ together with a bounded efficiency permits Carnot-type reasoning at the level of inference. Here $\Theta_{C}$ plays the formal role of a cold-reservoir temperature, limited by the representation noise. Unlike thermal engines, where efficiency measures work extraction and therefore involves a one minus structure, the efficiency here measures the fraction of attainable certainty achieved.
@@
Finally, the local efficiency bound $\Theta_{C}$ can also be shown to constrain global information efficiency. To see this, consider the ratio $\mathscr{I}/\mathscr{W}$, where $\mathscr{I}=\oint(\sigma^{2}/m\Theta)\,dm$ denotes the total information gained over a cycle and $\mathscr{W}=\oint(\sigma^{2}/m)\,dm$ represents the corresponding sampling work. This ratio can be seen as the weighted average of $\Theta^{-1}$ with respect to the measure $d\mathscr{W}=(\sigma^{2}/m)\,dm$ along the cycle. As such, it is bounded by the maximum and minimum values of $\Theta^{-1}$. Let $\Theta^{-1}$ attain its supremum along the cycle at $\Theta_{\inf}^{-1}$. It follows that $\mathscr{I}/\mathscr{W}=\langle\Theta^{-1}\rangle\leq\Theta_{\inf}^{-1}$. However, since $\Theta_{\inf}\geq 2m_{*}\sigma_{R}^{2}$, where $m_{*}$ is the minimum value of $m$ attained in the cycle, therefore $\mathscr{I}/\mathscr{W}\leq\Theta_{C}(m_{*})^{-1}$. That is, the global efficiency is bounded by the local efficiency evaluated at the point where $m$ reaches its minimum value along the cycle. In short, the representation noise sets the irreducible floor, while the sample size determines where the floor is evaluated. No choice of cycling through variance can overcome the constraint imposed by minimal sampling. Thus the global bound on information efficiency is encoded entirely by the local efficiency and by access to the smallest attainable sample size. This is analogous to Carnot’s result, which shows that efficiency depends only on the availability of extreme temperatures.
@@ Appendix B Appendix 2: Proof of the information inequality
$d\mathscr{I}=-\,\delta H_{\mathrm{relax}}=-\left(\frac{\partial H}{\partial m}\right)dm,$ (43)
where $H(\mu,m)$ is the entropy state function in (2).
@@
Let $\mu(t)$ be a cyclic stimulus that generates a closed stimulus–response trajectory in $(\mu,m)$ via the state space model described in Section 3. The change in information over a cycle obeys the second-law–type inequality

### v2

@@ 1 Introduction
The inference problem considered here departs from the study of static systems and instead addresses situations where the parameter of interest may change across operating conditions. We assume that inference is conducted independently over non-overlapping epochs, each treated as locally stationary in the statistical sense so that asymptotic estimation theory applies within each epoch. The sample size denotes the number of observations allocated to a single epoch and may vary between epochs to reflect changes in operating conditions. The variance may likewise change, reflecting shifts in the underlying parameter value, distribution, or the heterogeneity of the sampled environment. In this way, non-stationarity is accommodated through shifts in thermodynamic coordinates between epochs.
@@
The main contributions of the paper are as follows: (1) a thermodynamic state-space formulation of asymptotic inference in $(m,\sigma^{2})$ space, with an explicit uncertainty state function; (2) a reversed second-law–type cyclic inequality for mean inference and an integrating factor that yields a first-law–type balance relation; (3) a third-law–type lower bound on entropy imposed by representation noise, which sets a fundamental noise floor and limits efficiency; (4) optimal inference paths, global bounds on information gain, and a Carnot-like notion of inferential efficiency; and (5) a unification of familiar entropy–information identities, specifically the Gaussian-limit form of the de Bruijn and I–MMSE results, as coordinate projections of a single underlying structure.
@@ 3 Inference in sensory transduction and metrology
Three additional empirical assumptions are required to complete the description. First, the variance of stimulus fluctuations $\sigma^{2}(\mu)$ is assumed to be a continuous, non-decreasing function of the mean stimulus intensity $\mu$ (c.f. fluctuation scaling [4]), with $\mu$ varying between locally stationary epochs. Second, the sample size evolves dynamically, relaxing monotonically toward a unique equilibrium or optimal value $m_{eq}$. Third, since the variance increases monotonically with $\mu$, it is expected that the optimal sample size itself depends monotonically on stimulus magnitude $m_{eq}(\mu)$ although not necessarily with the same functional form as $\sigma^{2}(\mu)$. In what follows, sample size $m$ is treated as a continuous variable, with the approximation justified in the asymptotic regime.
@@
Sensory systems are not unique in exhibiting this statistical structure. A similar description arises whenever an unknown parameter is inferred from repeated, independent measurements. In metrology, one estimates a parameter $\theta$ of a distribution $p(x\mid\theta)$ from a set of finite samples $x_{1},\ldots,x_{m}$. In the asymptotic regime, efficient estimators become Gaussian with variance proportional to $1/m$, Fisher information becomes additive in $m$, and the estimator entropy admits a closed-form expression identical to (2) with $\sigma^{2}(\theta)=1/I(\theta)$ when representation noise is introduced. The inferential dynamics therefore run in the same direction as in sensory systems: increasing sample size reduces uncertainty. What metrology lacks, without further specification, is a dynamical sampling model (1), and the identification of uncertainty with a directly observable response. Moreover, without the analogue of $F=kH$, the thermodynamic structure describes limits on inferential efficiency rather than experimentally measurable results. Nevertheless, the underlying structure remains the same. In this sense, sensory transduction and measurement theory occupy the same position within a single inferential state space.
Next we turn to formulate a thermodynamics of inference with state variables $(m,\sigma^{2})$, entropy $H$, dynamics $\dot{m}=g(m,m_{eq})$, and trajectory-defined information production.
@@ 4 Foundations for a thermodynamic framework
where $\delta H_{\mathrm{flux}}=(\partial H/\partial\mu)\,d\mu$ and $\delta H_{\mathrm{relax}}=(\partial H/\partial m)\,dm$ are the coordinate contributions to the exact state differential $dH$.
Along relaxation trajectories, we define
$d\mathscr{I}=-\left(\frac{\partial H}{\partial m}\right)dm,$ (6)
or equivalently in rate form $\dot{\mathscr{I}}=-(\partial H/\partial m)\dot{m}$. We refer to this quantity as information production. Under the conditions discussed in Appendix 2, the theorem there gives the cyclic inequality
@@ 5 Identifying an integrating factor and a first law
The variance-driven contribution is
$dH_{\sigma}=\frac{d\sigma^{2}}{2(\sigma^{2}+m\sigma_{R}^{2})}$ (10)
and introducing the new state variable
@@
changes in entropy now takes the Clausius-like form $dH_{\sigma}=\Theta^{-1}d\sigma^{2}$. Here the state variable $\Theta$ plays a role similar to temperature as an integrating factor. As will be discussed later, at fixed large $m$, $\Theta$ is proportional to the inverse minimum mean square error (MMSE) for an averaged signal corrupted by additive Gaussian noise. An increase in variance produces a larger change in entropy when $m$ is small than when $m$ is large. For this reason, we refer to $\Theta$ as the uncertainty susceptibility.
It is important to emphasize that the two coordinates $m$ and $\sigma^{2}$ are treated independently and interpreted operationally. The quantity $m$ denotes the number of observations assigned to a single inference interval, with intervals taken to be statistically independent and non-overlapping. $\sigma^{2}$ can change across different epochs, reflecting changes in the underlying parameter value or in the statistical structure of the observations between epochs. Asymptotic relations apply within each interval, with no information carried between intervals. Accordingly, variations in $(m,\sigma^{2})$ represents shifts in operating conditions.
Interpreting variance in (10) as the quasi-heat is natural: variance is positive and quadratic, and often corresponds directly to signal power. Next, we introduce a first-law analogy directly from the state differential (9). Introducing $\Theta$, we get
@@
The quasi-work or sampling work term $(\sigma^{2}/m)dm$ can be interpreted as the incremental variance expenditure required to increase the effective sample size by $dm$. It has units of variance and reflects both the precision of the data and the effort associated with changing the sample size. When variance is constant, this term can be integrated between $m_{a}$ and $m_{b}$ to give $\sigma^{2}\log(m_{b}/m_{a})$. If $m_{a}=m_{b}$, the sampling work equals zero. This quantity has three main characteristics. First, it illustrates diminishing returns: relative gains arise when increasing samples by the same factor. Second, larger $\sigma^{2}$ (noisier data) requires more work to achieve the same improvement in precision. Third, it shows scale invariance, in that the work depends on the ratio $m_{b}/m_{a}$ rather than on the absolute difference between the sample sizes. In general, the integral of sampling work is path dependent since $(\sigma^{2}/m)\,dm$ is not an exact differential, and its value non-zero when evaluated over a cycle. To illustrate this, consider a system with variance $\sigma_{a}^{2}$ sampled with sample size $m_{a}$. At $t=0$, the system is perturbed so that the variance increases to $\sigma_{b}^{2}\geq\sigma_{a}^{2}$ while the sample size is increased to $m_{b}$. At a later time, the variance returns to $\sigma_{a}^{2}$ and the sample size returns to $m_{a}$. The sampling work over this cycle is
@@
$=(\sigma_{b}^{2}-\sigma_{a}^{2})\log\!\left(\frac{m_{b}}{m_{a}}\right),$ (20)
which can be interpreted geometrically as the weighted area enclosed in $(m,\sigma^{2})$ space.
@@ 6 Other thermodynamic relationships, and a third law
$\left(\frac{\partial H}{\partial\sigma^{2}}\right)_{m}$ $=\frac{1}{\Theta},\quad$ $\left(\frac{\partial H}{\partial m}\right)_{\sigma^{2}}$ $=-\frac{\sigma^{2}}{m\Theta},$ (22)
@@ 7.1 Information gain
$\Delta\mathscr{I}=-\displaystyle\int_{m_{a}}^{m_{b}}\left(\frac{\partial H}{\partial m}\right)_{\sigma^{2}}\,dm=\displaystyle\int_{m_{a}}^{m_{b}}\frac{1}{\Theta}\frac{\sigma^{2}}{m}\,dm,$ (23)
@@ 7.2 Information efficiency
The estimator variance is $\sigma^{2}/m$, and so increasing the sample size reduces the effective variance. Two complementary information identities emerge depending on which coordinate is treated as the control variable. When the variance is varied at fixed sample size, (29) corresponds to the Gaussian form of the I–MMSE relation [6]. For an additive Gaussian channel with signal-to-noise ratio $\mathrm{SNR}=\sigma^{2}/(m\sigma_{R}^{2})$, the minimum mean-square error satisfies
@@
Alternatively, if the sample size is varied while the variance is held fixed, entropy changes from (22) vary as
$\left(\frac{\partial H}{\partial m}\right)_{\sigma^{2}}=-\frac{\sigma^{2}}{m\Theta}.$ (31)
In this sampling-driven setting, the entropy evolution corresponds to the Gaussian form of de Bruijn’s identity [3]. Increasing $m$ reduces the effective noise variance $\sigma^{2}/m$, producing a diffusion process in reverse time.
The appearance of both de Bruijn’s identity and the I–MMSE relation reflects the Gaussian structure underlying the entropy function. More importantly, they arise as complementary coordinate projections of a single underlying thermodynamic structure. This unification is encoded by the state function $\Theta=2(\sigma^{2}+m\sigma_{R}^{2})$, which governs joint sampling- and variance-driven information transduction.
@@
where normalization by the input signal variance $\sigma^{2}/m$ defines a dimensionless quantity $\eta$ characterizing local information efficiency. The efficiency obeys $0\leq\eta\leq 1$, with $\eta=1$ attained only when $m\to\infty$ which is impossible since $m$ is a resource. The existence of a temperature-like variable $\Theta$ together with a bounded efficiency permits Carnot-type reasoning at the level of inference. Here $\Theta_{C}$ plays the formal role of a cold-reservoir temperature, limited by the representation noise. Unlike thermal engines, where efficiency measures work extraction and therefore involves a one minus structure, the efficiency here measures the fraction of attainable certainty achieved.
@@
Finally, the measure of local efficiency can also be shown to constrain global information efficiency. To see this, consider the ratio $\mathscr{I}/\mathscr{W}$, where $\mathscr{I}=\oint(\sigma^{2}/m\Theta)\,dm$ denotes the total information gained over a cycle and $\mathscr{W}=\oint(\sigma^{2}/m)\,dm$ represents the corresponding sampling work. This ratio can be interpreted as a weighted average of $\Theta^{-1}$ with respect to the measure $d\mathscr{W}=(\sigma^{2}/m)\,dm$ along the cycle. Consequently, it must lie between the minimum and maximum values of $\Theta^{-1}$ attained along the path. Let $\Theta^{-1}$ attain its supremum along the cycle at $\Theta_{\inf}^{-1}$. It follows that $\mathscr{I}/\mathscr{W}=\langle\Theta^{-1}\rangle\leq\Theta_{\inf}^{-1}$. From the definition $\Theta=2(\sigma^{2}+m\sigma_{R}^{2})$, it follows that $\Theta_{\inf}\geq 2m_{*}\sigma_{R}^{2}$, where $m_{*}$ is the minimum value of $m$ attained along the cycle. Therefore $\mathscr{I}/\mathscr{W}\leq\Theta_{C}(m_{*})^{-1}$. Thus the global information efficiency is bounded by the local efficiency evaluated at the point where the sample size reaches its minimum along the cycle. In other words, the representation noise $\sigma_{R}^{2}$ sets the irreducible floor, while the smallest attainable sample size determines where this floor is realized. In summary, global efficiency is limited by the extreme values of $\Theta$ encountered along the path. In particular, the upper bound depends on access to the smallest achievable sample size. This mirrors Carnot’s result in thermodynamics, where the efficiency of a cycle depends only on the availability of extreme temperatures.
@@ Appendix B Appendix 2: Proof of the information inequality
$d\mathscr{I}=-\left(\frac{\partial H}{\partial m}\right)dm,$ (44)
where $H$ is the entropy state function in (2) and $dm$ is evaluated along trajectories generated by $\dot{m}=g(m,m_{eq})$. Since $\sigma^{2}$ is a monotonic function of the stimulus parameter $\mu$, we will refer to the state space using coordinates $(\mu,m)$ for the purposes of analyzing stimulus–response cycles.
@@
Let $\mu(t)$ be a cyclic stimulus that generates a closed stimulus–response trajectory in $(\mu,m)$ via the dynamical state-space model described in Section 3. The change in information over a cycle obeys the second-law–type inequality